# bibliography.bib

### v1

@@ Abstract
$\mathfrak{hc}^{\prime}(\wedge^{n+1}\mathcal{S}_{n})\star\hat{\mathcal{T}}_{w_{0}}\simeq\wedge^{n+1}V\otimes_{\overline{\mathbf{Q}}_{\ell}}\hat{\mathcal{T}}_{w_{0}}=0.$
This implies that the Steinberg character sheaf $L_{St}$ can not be a direct summand of $\mathfrak{hc}^{\prime}(\wedge^{n+1}\mathcal{S}_{n})$. Indeed, for any irreducible unipotent character sheaf $L\neq L_{St}$, we have $\mathfrak{hc}^{\prime}(L)\star\hat{\mathcal{T}}_{w_{0}}=0$ by 3.4.7 (c) and [lusztigTruncatedConvolutionCharacter2014][Proposition 2.7]. However, the category of unipotent character sheaves is not sent to 0 by the functor $\mathfrak{hc}^{\prime}(-)\star\hat{\mathcal{T}}_{w_{0}}$, we get that $\mathfrak{hc}^{\prime}(L_{St})\star\hat{\mathcal{T}}_{w_{0}}\neq 0$. This concludes the proof. ∎
3.5   $\operatorname{Rep}(\mathbf{G})^{+}$ and $\operatorname{Rep}(\mathbf{G})$-action.
@@
We will use the framework of [deligneCategorieRepresentationsGroupe2007][Section 10]. Let ${\operatorname{Rep}^{+}(\mathbf{GL}(t=n),\mathbf{K})}$ stand for the full Caroubian subcategory of the Deligne category $\operatorname{Rep}(\mathbf{GL}(t=n),\mathbf{K})$ on objects $(U,\varnothing)$ in the notations of loc. cit. Here notation $(t=n)$ means that we specialize the indeterminate $t$ in the definition of $\operatorname{Rep}(\mathbf{GL}(t),\mathbf{K})$ to $n$. The category ${\operatorname{Rep}^{+}(\mathbf{GL}(t=n),\mathbf{K})}$ is just the idempotent completion of a category with objects $[m]\in\mathbb{N}$ and morphisms given by 1-dimensional unoriented flat bordisms. There is a functor ${\operatorname{Rep}^{+}(\mathbf{GL}(t=n),\mathbf{K})}\to\mathcal{C}$, sending the generating object $[1]$ to $X$. The fact that $\wedge^{n+1}X=0$ implies, by the second fundamental theorem of invariant theory of the general linear group (see, for example, [procesiLieGroupsApproach2007][Theorem 6.1], that this functor factors through $\operatorname{Rep}(\mathbf{G})^{+}$. This proves (a).
We now prove (b). For $W\in\operatorname{Rep}(\mathbf{G})^{+}$ let $\phi_{W}:\mathcal{M}\to\mathcal{M}$ be the corresponding functor. Let $\det,V\in\operatorname{Rep}(\mathbf{G})^{+}$ be the determinant and standard representations, respectively. Define $\phi_{V^{\vee}}=\phi_{\wedge^{n-1}V}\circ\phi_{\det}^{-1}$. It is easy to see that this defines an action of ${\operatorname{Rep}(\mathbf{GL}(t=n),\mathbf{K})},\operatorname{Rep}(\mathbf{G})$ as in (a). ∎
@@
By Theorem 3.4.1 and Lemma 3.5.1 (a), there is a functor $\zeta:\operatorname{Rep}(\mathbf{G})^{+}\to P_{G}(G)$. Projecting to the category of unipotent character sheaves, we have a functor $\hat{\zeta}:\operatorname{Rep}(\mathbf{G})^{+}\to\operatorname{pro}P^{\mathbf{1}}(G),V\mapsto\left(\zeta(V)\star E_{k}^{0}\right)_{k=0}^{\infty}$ (note that the projections of $\wedge^{k}\mathcal{S}_{n}$ are perverse by [hhh][Lemma 5.6.3]). By Proposition 3.4.7 (b), (d), $P^{\mathbf{1}}(G)$ acts on $\mathcal{H}_{n}^{\mathrm{perf}}$. Moreover, the action of $\det\in\operatorname{Rep}(\mathbf{G})^{+}$ obtained is by $\mathfrak{hc}(\mathcal{S}_{\wedge^{n}})\star\hat{\Delta}_{e}\simeq\hat{\Delta}_{w_{0}}^{\star 2}$, where the last isomorphism is [hhh][Corollary 5.7.6]. From Proposition 2.1.1 (d) it follows that this action is by an invertible functor, and so by Lemma 3.5.1 (b) we get the result. ∎
Let us retain the notation from the proof of Corollary 3.5.2: for $V\in\operatorname{Rep}(\mathbf{G})^{+}$ let $\hat{\zeta}(V)\in\operatorname{pro}P_{G}(G)$ be the corresponding unipotent character (pro-)sheaf. For an arbitrary $V\in\operatorname{Rep}(\mathbf{G})$, let $\phi_{V}:\mathcal{H}_{n}^{\mathrm{perf}}\to\mathcal{H}_{n}^{\mathrm{perf}}$ be the corresponding functor. We have $\phi_{V}(-)=\mathfrak{hc}(\hat{\zeta}(V))\star-$ for $V\in\operatorname{Rep}(\mathbf{G})^{+}$.
@@
The main result of this section is
@@
    $\varpi(\mathcal{O}_{\mathbf{St}_{\mathfrak{g}}}(\lambda,\mu))=\mathfrak{J}_{\lambda}\star\hat{\mathcal{T}}_{w_{0}}\star\mathfrak{J}_{\mu}$ for any two weights $\lambda,\mu\in\operatorname{Rep}(\mathbf{T})$.
@@
    $\varpi(V\otimes\mathcal{O}_{\mathbf{St}_{\mathfrak{g}}})=\phi_{V}\hat{\mathcal{T}}_{w_{0}}$ for any representation $V$ of $G$.
Proof.
We employ the general construction of [bezrukavnikovTwoGeometricRealizations2016]. Let $\mathfrak{t}$ be the Lie algebra of $\mathbf{T}$.
Let $\mathcal{B}=\mathbf{G}/\mathbf{B},\mathbf{Y}=\mathbf{G}/\mathbf{U}$.
Let $\mathbf{C}_{\tilde{\mathfrak{g}}}$ be the preimage in $\mathfrak{g}\times\mathbf{Y}$ of $\tilde{\mathfrak{g}}$ under the natural projection ${\mathfrak{g}\times\mathbf{Y}\to\mathfrak{g}\times\mathcal{B}}$. Since the right $\mathbf{T}$-action on $\mathbf{Y}$ is free, we have $\mathrm{coh}_{\mathbf{G}\times\mathbf{T}}(\mathbf{C}_{\tilde{\mathfrak{g}}})=\mathrm{coh}_{G}(\tilde{\mathfrak{g}})$.
Let $\overline{\mathbf{Y}}=\operatorname{Spec}\Gamma(\mathcal{O}_{\mathbf{Y}})$ be the affine closure of $\mathbf{Y}$.
Now pick a representation $\mathbf{V}$ of $\mathbf{G}$, such that
@@
    $\mathbf{V}$ is a multiplicity-free sum of irreducible representations;
@@
    $\mathbf{Y}$ is a locally closed subvariety of $\mathbf{V}$, namely an orbit of a highest-weight vector;
@@
    $\overline{\mathbf{Y}}$ is a closure of $\mathbf{Y}$ inside $\mathbf{V}$.
Choose an action of $\mathfrak{t}$ on $\mathbf{V}$ so that $t\in\mathfrak{t}$ acts on an irreducible summand with the highest weight $\lambda$ by a constant $\langle\lambda,t\rangle$.
Let $\overline{\mathbf{C}}_{\tilde{\mathfrak{g}}}$ be a closed subscheme of $\mathfrak{g}\times\mathfrak{t}\times\overline{\mathbf{Y}}$ given by
$\overline{\mathbf{C}}_{\tilde{\mathfrak{g}}}=\{(x,t,v)\in\mathfrak{g}\times\mathfrak{t}\times\overline{\mathbf{Y}}\subset\mathfrak{g}\times\mathfrak{t}\times\mathbf{V}:x(v)=t(v)\}.$
$\mathbf{C}_{\tilde{\mathfrak{g}}}$ becomes an open subset in $\overline{\mathbf{C}}_{\tilde{\mathfrak{g}}}$, given by $(x,t,v)$ with $v$ having a non-zero projection to all irreducible summands of $\mathbf{V}$.
Let $\overline{\mathbf{C}}_{\textrm{St}}$ be the preimage of the diagonal under the map $\overline{\mathbf{C}}_{\tilde{\mathfrak{g}}}\times\overline{\mathbf{C}}_{\tilde{\mathfrak{g}}}\to\mathfrak{g}\times\mathfrak{g}$. We have an open subset $\mathbf{C}_{\mathbf{St}_{\mathfrak{g}}}\subset\overline{\mathbf{C}}_{\mathbf{St}_{\mathfrak{g}}}$ with a free $\mathbf{T}^{2}$-action, and $\mathbf{C}_{\mathbf{St}_{\mathfrak{g}}}/\mathbf{T}^{2}=\mathbf{St}_{\mathfrak{g}}.$
For a variety $\mathbf{X}$ with an action of an algebraic group $\mathbf{H}$, let $\mathrm{coh}^{\mathrm{fr}}_{\mathbf{H}}(\mathbf{X})$ be the full subcategory of $\mathrm{coh}_{\mathbf{H}}(\mathbf{X})$, consisting of objects of the form $V\otimes\mathcal{O}_{\mathbf{X}}$, $V\in\operatorname{Rep}(\mathbf{H})$.
For a dominant weight $\lambda$ let $\mathbf{V}_{\lambda}\in\operatorname{Rep}(\mathbf{G})$ be the irreducible representation with the highest weight $\lambda$. We have the following
Proposition 4.1.2 ([bezrukavnikovTwoGeometricRealizations2016], Corollary 18 and Section 4.4.1).
@@
    An action of $\mathcal{O}_{\mathfrak{t}}$ on $F$ by endomorpisms, so that for $f\in\mathcal{O}_{\mathfrak{t}}$ we have $f_{\mathbf{V}_{1}\otimes\mathbf{V}_{2}}=f_{\mathbf{V}_{1}}\otimes Id_{\mathbf{V}_{2}}=Id_{\mathbf{V}_{1}}\otimes f_{\mathbf{V}_{2}}$.
@@
    A “highest weight arrow” $w_{\lambda}:F(\lambda)\to F(\mathbf{V}_{\lambda})$ making the following diagrams commutative:
@@
    ${F({\lambda})}$${F(\mathbf{V}_{\lambda})}$${F({\lambda})}$${F(\mathbf{V}_{\lambda})}$$\scriptstyle{w_{\lambda}}$$\scriptstyle{\lambda}$$\scriptstyle{E_{\lambda}}$$\scriptstyle{w_{\lambda}}$
    where the right vertical map is the action of the element $\lambda\in\mathfrak{t}^{\vee}\subset\mathcal{O}(\mathfrak{t})$ coming from (c).
Then the tensor functor $F$ extends uniquely to an action of $\mathrm{coh}^{\mathrm{fr}}_{\mathbf{G}\times\mathbf{T}}(\overline{\mathbf{C}}_{\tilde{\mathfrak{g}}})$ on $\mathcal{\mathbf{C}}$. If we have two copies of the above data, we get an action of $\mathrm{coh}^{\mathrm{fr}}_{\mathbf{G}^{2}\times\mathbf{T}^{2}}(\overline{\mathbf{C}}_{\tilde{\mathfrak{g}}}\times\overline{\mathbf{C}}_{\tilde{\mathfrak{g}}})$. In this case, assume, moreover, that
@@
    the action of $\operatorname{Rep}(\mathbf{G}\times\mathbf{G})$ factors through the diagonal $\mathbf{G}\to\mathbf{G}\times\mathbf{G}$.
@@
    this action of $\operatorname{Rep}(\mathbf{G})$ is by product with central objects in a braided category $\mathcal{C}$, with endomorphism $E$ from (b) respecting the central structure.
@@
First, the action of $\operatorname{Rep}(\mathbf{G}\times\mathbf{G})$ is given by Corollary 3.5.2 applied to left and right action of $P^{\mathbf{1}}(G)$ on $\mathcal{H}_{n}^{\mathrm{perf}}$. The action of $\operatorname{Rep}(\mathbf{T}\times\mathbf{T})$ is given by left and right convolutions with Jucys-Murphy sheaves $\mathfrak{J}_{\lambda}$.
Action of $f\otimes g\in\mathcal{O}_{\mathfrak{t}\times\mathfrak{t}}$ is given by the action of $f$ and $g$ coming from left and right monodromy, respectively.
@@
Let $V_{n},\mathbf{V}_{n}$ stand for the standard $n$-dimensional representation of $\textrm{GL}_{n},\mathbf{GL}_{n}$ with fixed bases $(e_{1},\dots,e_{n}),(\mathbf{e}_{1},\dots,\mathbf{e}_{n})$, respectively. For $k<n$ consider the subspace $V_{k}=\operatorname{span}_{\mathbf{k}}(e_{1},\dots,e_{k})$, respectively $\mathbf{V}_{k}=\operatorname{span}_{\mathbf{K}}(\mathbf{e}_{1},\dots,\mathbf{e}_{k})$ as a standard representation of $\textrm{GL}_{k}=\textrm{GL}(V_{k}),\mathbf{GL}_{k}=\mathbf{GL}(\mathbf{V}_{k})$. Let $\operatorname{triv}_{\mathbf{GL}_{n-k}}$ stand for the trivial representation of $\mathbf{GL}_{n-k}$. Consider
=${}_{k}=\hat{\zeta}(\wedge^{k}\mathbf{V}_{k})\boxtimes\hat{\zeta}(\operatorname{triv}_{\mathbf{GL}_{n-k}}),$
@@
We have $\hat{\zeta}(\wedge^{k}\mathbf{V}_{n})=\operatorname{Ind}_{P_{k}}^{G}({}_{k})$. For a an arbitrary reductive group $G$, let $G^{rss}$ stand for the open subset of regular semisimple elements. k is the IC-extension of the (pro-)local system on $L_{k}^{rss}=\textrm{GL}_{k}^{rss}\times\textrm{GL}_{n-k}^{rss}$. It will be convenient to describe this (pro-)local system explicitly. Recall that the category of pro-unipotent local systems on $T$, as defined in [bezrukavnikovKoszulDualityKacMoody2014], can be identified with the category of $\mathbf{K}[[\mathbf{V}_{n}]]$-modules. Let
@@
$\mathfrak{hc}\hat{\zeta}(\wedge^{k}\mathbf{V}_{n})=\mathfrak{hc}_{B}^{P_{k}}\mathfrak{hc}_{P_{k}}^{G}\chi_{P_{k}}^{G}\iota_{P_{k}}({}_{k}).$
Let $\omega_{k}$ be the highest weight of $\wedge^{k}V_{n}$. We have
$\mathfrak{hc}_{B}^{P_{k}}\circ\iota_{P_{k}}({}_{k})\simeq\mathfrak{J}_{\omega_{k}}=\mathfrak{J}_{1}\star\dots\star\mathfrak{J}_{k},$
see the [hhh][Lemma 6.2.4]. We also have the natural morphism
$r_{k}:\iota_{P_{k}}({}_{k})\to\mathfrak{hc}_{P_{k}}^{G}\chi_{P_{k}}^{G}\iota_{P_{k}}({}_{k}),$
see Subsection 3.3.1.
@@
[…]
Here $w_{\omega_{k}}=\mathfrak{hc}_{B}^{P_{k}}(r_{k}),E_{k}=\mathfrak{hc}(\varepsilon_{k})$, and $\omega_{k}$ comes from a monodromy action on $\mathfrak{J}_{\omega_{k}}$.
We define $w_{\omega_{k}}$ to be the highest weight arrows for the images of fundamental representations $\wedge^{k}\mathbf{V}_{n}$. Note that, since the highest weight vector of any representation can be given by a product of the highest weight vectors of fundamental representations in their tensor product, we may define, for a highest weight $\lambda=\sum a_{i}\omega_{i},w_{\lambda}=\prod_{i}w_{\omega_{i}}^{\star a_{i}}$.
@@
$E_{\omega_{k}}$ obtain from $E_{1}$ by imposing condition (b) is equal to $E_{k}$ defined above.
@@
$\operatorname{Ind}^{G}_{P_{1}}\operatorname{Res}^{G}_{P_{1}}(\cdot)\to\mathcal{S}_{n}\star(\cdot),$
@@
$\left(\operatorname{Ind}^{G}_{P_{1}}\operatorname{Res}^{G}_{P_{1}}\right)^{\circ k}(\cdot)\to\mathcal{S}_{n}^{\star k}\star(\cdot).$
By the proof of Proposition 3.4.2, it gives an isomorphism on the direct summands corresponding to the exterior power. If a character sheaf $\mathcal{F}$ is given by the IC-extension of a local system corresponding to the $A_{n}$-module $M$, then $\operatorname{Ind}^{G}_{P}\operatorname{Res}^{G}_{P}\mathcal{F}$ is given by the IC-extension corresponding to a $A_{n}$-module $\mathbf{K}[S_{n}]\otimes_{\mathbf{K}[S_{n-1}]}M$.
Consider an action of $S_{n}$ on $\mathbf{V}_{n}$ by permutation of the basis $(\mathbf{e}_{i})$. Choosing an isomorphism of $S_{n}$-modules
@@
and comparing the actions under this identification, we get that the action of $E_{\omega_{k}}$ on the corresponding direct summand is the same as given by $E_{k}$. ∎
4.3   Filtration by Jucys-Murphy sheaves.

### v2

@@ Abstract
$\mathfrak{hc}^{\prime}(\wedge^{n+1}\mathcal{S}_{n})\star\hat{\mathcal{T}}_{w_{0}}\simeq\wedge^{n+1}\mathbf{V}\otimes_{\overline{\mathbf{Q}}_{\ell}}\hat{\mathcal{T}}_{w_{0}}=0.$
Note that for any irreducible unipotent character sheaf $L\neq L_{St}$, we have $\mathfrak{hc}^{\prime}(L)\star\hat{\mathcal{T}}_{w_{0}}=0$ by 3.4.7 (c) and [Lus15, Proposition 2.7]. However, the category of unipotent character sheaves is not sent to 0 by the functor $\mathfrak{hc}^{\prime}(-)\star\hat{\mathcal{T}}_{w_{0}}$, and so we get that $\mathfrak{hc}^{\prime}(L_{St})\star\hat{\mathcal{T}}_{w_{0}}\neq 0$. This implies that the Steinberg character sheaf $L_{St}$ can not be a direct summand of $\mathfrak{hc}^{\prime}(\wedge^{n+1}\mathcal{S}_{n})$. This concludes the proof. ∎
3.5  The $\operatorname{Rep}(\mathbf{G})$-action.
@@
We will use the framework of [Del07, Section 10]. Let $\operatorname{Rep}^{+}(\mathbf{GL}(t=n),\mathbf{K})$} stand for the full idempotent-complete subcategory of the Deligne category $\operatorname{Rep}(\mathbf{GL}(t=n),\mathbf{K})$ with objects $(U,\varnothing)$ in the notations of ibid. Here notation $t=n$ means that we specialize the indeterminate $t$ in the definition of $\operatorname{Rep}(\mathbf{GL}(t),\mathbf{K})$ to $n$. The category $\operatorname{Rep}^{+}(\mathbf{GL}(t=n),\mathbf{K})$} is just the idempotent completion of a category with objects $m]\in\mathbb{N}$ and morphisms given by 1-dimensional unoriented flat bordisms. There is a functor ${\operatorname{Rep}^{+}(\mathbf{GL}(t=n),\mathbf{K})}\to\mathcal{C}$, sending the generating object $[1]$ to $X$. The fact that $\wedge^{n+1}X=0$ implies, by the second fundamental theorem of invariant theory of the general linear group (see, for example, [Pro07, Theorem 6.1], that this functor factors through $\operatorname{Rep}(\mathbf{G})^{+}$. This proves (a).
We now prove (b). For $\mathbf{W}\in\operatorname{Rep}(\mathbf{G})^{+}$, let $\phi_{\mathbf{W}}:\mathcal{M}\to\mathcal{M}$ be the corresponding functor. Let $\mathbf{det},\mathbf{V}\in\operatorname{Rep}(\mathbf{G})^{+}$ be the determinant and standard representations, respectively. Define $\phi_{\mathbf{V}^{\vee}}=\phi_{\wedge^{n-1}\mathbf{V}}\circ\phi_{\mathbf{det}}^{-1}$. It is easy to see that this defines an action of ${\operatorname{Rep}(\mathbf{GL}(t=n),\mathbf{K})},\operatorname{Rep}(\mathbf{G})$ as in (a). ∎
@@
By Theorem 3.4.1 and Lemma 3.5.1 (a), there is a functor $\zeta:\operatorname{Rep}(\mathbf{G})^{+}\to P_{G}(G)$. Projecting to the category of unipotent character sheaves, we have a functor $\hat{\zeta}:\operatorname{Rep}(\mathbf{G})^{+}\to\operatorname{pro}P^{\mathbf{1}}(G)$, sending a representation $\mathbf{W}$ to the system $\left(\zeta(\mathbf{W})\star E_{k}^{0}\right)_{k=0}^{\infty}$ (note that the projections of $\wedge^{k}\mathcal{S}_{n}$ are perverse by [BT22, Lemma 5.6.3]). By Proposition 3.4.7 (b), (d), $P^{\mathbf{1}}(G)$ acts on $\mathcal{H}_{n}^{\mathrm{perf}}$. Moreover, the action of $\mathbf{det}\in\operatorname{Rep}(\mathbf{G})^{+}$ obtained is by $\mathfrak{hc}(\mathcal{S}_{\wedge^{n}})\star\hat{\Delta}_{e}\simeq\hat{\Delta}_{w_{0}}^{\star 2}$, where the last isomorphism is [BT22, Corollary 5.7.6]. From Proposition 2.1.1 (d) it follows that this action is by an invertible functor, and so by Lemma 3.5.1 (b) we get the result. ∎
Let us retain the notation from the proof of Corollary 3.5.2: for $\mathbf{W}\in\operatorname{Rep}(\mathbf{G})^{+}$ let $\hat{\zeta}(\mathbf{W})\in\operatorname{pro}P_{G}(G)$ be the corresponding unipotent character (pro-)sheaf. For an arbitrary $\mathbf{W}\in\operatorname{Rep}(\mathbf{G})$, let $\phi_{\mathbf{W}}:\mathcal{H}_{n}^{\mathrm{perf}}\to\mathcal{H}_{n}^{\mathrm{perf}}$ be the corresponding functor. We have $\phi_{\mathbf{W}}(-)\simeq\mathfrak{hc}(\hat{\zeta}(\mathbf{W}))\star$- for $\mathbf{W}\in\operatorname{Rep}(\mathbf{G})^{+}$.
@@
This section is occupied with the proof of the following
@@
    $\varpi(\mathcal{O}_{\mathbf{St}_{\mathfrak{g}}}(\boldsymbol{\lambda},\boldsymbol{\mu}))=\mathfrak{J}_{\boldsymbol{\lambda}}\star\hat{\mathcal{T}}_{w_{0}}\star\mathfrak{J}_{\boldsymbol{\mu}}$ for any two weights $\boldsymbol{\lambda},\boldsymbol{\mu}\in\operatorname{Rep}(\mathbf{T})$.
@@
    $\varpi(\mathbf{W}\otimes\mathcal{O}_{\mathbf{St}_{\mathfrak{g}}})=\phi_{\mathbf{W}}\hat{\mathcal{T}}_{w_{0}}$ for any representation $\mathbf{W}$ of $G$.
[…]
Now pick a representation $\mathbf{V_{0}}$ of $\mathbf{G}$, such that
@@
    $\mathbf{V_{0}}$ is a multiplicity-free sum of irreducible representations;
@@
    $\mathbf{Y}$ is a locally closed subvariety of $\mathbf{V_{0}}$, namely an orbit of a highest-weight vector;
@@
    $\overline{\mathbf{Y}}$ is the closure of $\mathbf{Y}$ inside $\mathbf{V_{0}}$.
Choose an action of $\mathfrak{t}$ on $\mathbf{V_{0}}$ so that $\mathbf{t}\in\mathfrak{t}$ acts on an irreducible summand with the highest weight $\boldsymbol{\lambda}$ by the constant $\langle\boldsymbol{\lambda},\mathbf{t}\rangle$. Let $\overline{\mathbf{C}}_{\tilde{\mathfrak{g}}}$ be the closed subscheme of $\mathfrak{g}\times\mathfrak{t}\times\overline{\mathbf{Y}}$ given by
$\overline{\mathbf{C}}_{\tilde{\mathfrak{g}}}=\{(\mathbf{x},\mathbf{t},\mathbf{v})\in\mathfrak{g}\times\mathfrak{t}\times\overline{\mathbf{Y}}\subset\mathfrak{g}\times\mathfrak{t}\times\mathbf{V_{0}}:\mathbf{x}(\mathbf{v})=\mathbf{t}(\mathbf{v})\}.$
$\mathbf{C}_{\tilde{\mathfrak{g}}}$ becomes an open subset in $\overline{\mathbf{C}}_{\tilde{\mathfrak{g}}}$, given by $(\mathbf{x},\mathbf{t},\mathbf{v})$ with $\mathbf{v}$ having a non-zero projection to all irreducible summands of $\mathbf{V_{0}}$. Let $\overline{\mathbf{C}}_{\textrm{St}}$ be the preimage of the diagonal under the map $\overline{\mathbf{C}}_{\tilde{\mathfrak{g}}}\times\overline{\mathbf{C}}_{\tilde{\mathfrak{g}}}\to\mathfrak{g}\times\mathfrak{g}$. We have an open subset $\mathbf{C}_{\mathbf{St}_{\mathfrak{g}}}\subset\overline{\mathbf{C}}_{\mathbf{St}_{\mathfrak{g}}}$ with a free $\mathbf{T}^{2}$-action, coming from the right $\mathbf{T}$-action on $\mathbf{Y}=\mathbf{G}/\mathbf{U}$, and an isomorphism $\mathbf{C}_{\mathbf{St}_{\mathfrak{g}}}/\mathbf{T}^{2}=\mathbf{St}_{\mathfrak{g}}.$
For a variety $\mathbf{X}$ with an action of an algebraic group $\mathbf{H}$, let $\mathrm{coh}^{\mathrm{fr}}_{\mathbf{H}}(\mathbf{X})$ be the full subcategory of $\mathrm{coh}_{\mathbf{H}}(\mathbf{X})$, consisting of objects of the form $\mathbf{W}\otimes\mathcal{O}_{\mathbf{X}}$ with $\mathbf{W}\in\operatorname{Rep}(\mathbf{H})$.
For a dominant weight $\boldsymbol{\lambda}$ let $\mathbf{V}_{\boldsymbol{\lambda}}\in\operatorname{Rep}(\mathbf{G})$ be the irreducible representation with the highest weight $\boldsymbol{\lambda}$. The following statement is [Bez16, Corollary 18 and Section 4.4.1].
Proposition 4.1.2.
@@
    An action of $\mathcal{O}_{\mathfrak{t}}$ on $F$ by endomorphisms, such that for $f\in\mathcal{O}_{\mathfrak{t}}$ we have $f_{\mathbf{V}_{1}\otimes\mathbf{V}_{2}}=f_{\mathbf{V}_{1}}\otimes Id_{F(\mathbf{V}_{2})}=Id_{F(\mathbf{V}_{1})}\otimes f_{\mathbf{V}_{2}}$.
@@
    A “highest weight arrow” $w_{{\boldsymbol{\lambda}}}:F({\boldsymbol{\lambda}})\to F(\mathbf{V}_{\boldsymbol{\lambda}})$ making the following diagrams commutative:
    $F({{\boldsymbol{\lambda}}})\otimes_{\mathcal{C}}F({{\boldsymbol{\mu}}})$$F\left({\boldsymbol{\lambda}}+{\boldsymbol{\mu}}\right)$$F(\mathbf{V}_{\boldsymbol{\lambda}})\otimes_{\mathcal{C}}F(\mathbf{V}_{\boldsymbol{\mu}})\simeq F(\mathbf{V}_{\boldsymbol{\lambda}}\otimes\mathbf{V}_{\boldsymbol{\mu}})$$F(\mathbf{V}_{{\boldsymbol{\lambda}}+{\boldsymbol{\mu}}})$$\scriptstyle{w_{\boldsymbol{\lambda}}\otimes_{\mathcal{C}}w_{\boldsymbol{\mu}}}$$\scriptstyle{w_{{\boldsymbol{\lambda}}+{\boldsymbol{\mu}}}}$ (9)
    $F({{\boldsymbol{\lambda}}})$$F(\mathbf{V}_{\boldsymbol{\lambda}})$$F({{\boldsymbol{\lambda}}})$$F(\mathbf{V}_{\boldsymbol{\lambda}})$$\scriptstyle{w_{{\boldsymbol{\lambda}}}}$$\scriptstyle{{\boldsymbol{\lambda}}}$$\scriptstyle{E_{\mathbf{V}_{\boldsymbol{\lambda}}}}$$\scriptstyle{w_{{\boldsymbol{\lambda}}}}$
    where the right vertical map is the action of the element ${\boldsymbol{\lambda}}\in\mathfrak{t}^{\vee}\subset\mathcal{O}(\mathfrak{t})$ coming from (c).
@@
    the action of $\operatorname{Rep}(\mathbf{G}\times\mathbf{G})$ factors through the action of $\operatorname{Rep}(\mathbf{G})$ using the restriction along the diagonal embedding $\mathbf{G}\to\mathbf{G}\times\mathbf{G}$.
@@
    this action of $\operatorname{Rep}(\mathbf{G})$ is by product with central objects in $\mathcal{C}$, with endomorphism $E$ from (b) respecting the central structure.
@@
First, the action of $\operatorname{Rep}(\mathbf{G}\times\mathbf{G})$ is given by Corollary 3.5.2 applied to left and right action of $P^{\mathbf{1}}(G)$ on $\mathcal{H}_{n}^{\mathrm{perf}}$. The action of $\operatorname{Rep}(\mathbf{T}\times\mathbf{T})$ is given by left and right convolutions with Jucys–Murphy sheaves $\mathfrak{J}_{\boldsymbol{\lambda}}$.
The action of $f\otimes g\in\mathcal{O}_{\mathfrak{t}\times\mathfrak{t}}$ is given by the action of $f$ and $g$ coming from left and right monodromy, respectively.
@@
Let $V_{n},\mathbf{V}_{n}$ stand for the standard $n$-dimensional representation of $\textrm{GL}_{n},\mathbf{GL}_{n}$ with fixed bases $(e_{1},\dots,e_{n}),(\mathbf{e}_{1},\dots,\mathbf{e}_{n})$, respectively. For $k<n$ consider the subspace $V_{k}=\operatorname{span}_{\mathbf{k}}(e_{1},\dots,e_{k})$, respectively $\mathbf{V}_{k}=\operatorname{span}_{\mathbf{K}}(\mathbf{e}_{1},\dots,\mathbf{e}_{k})$, as a standard representation of $\textrm{GL}_{k}=\textrm{GL}(V_{k}),\mathbf{GL}_{k}=\mathbf{GL}(\mathbf{V}_{k})$. Let $\operatorname{triv}_{\mathbf{GL}_{n-k}}$ stand for the trivial representation of $\mathbf{GL}_{n-k}$. Consider
$_{k}=\hat{\zeta}(\wedge^{k}\mathbf{V}_{k})\boxtimes\hat{\zeta}(\operatorname{triv}_{\mathbf{GL}_{n-k}}),$
@@
[…]
@@
[…]
Let $\boldsymbol{\omega}_{k}$ be the highest weight of $\wedge^{k}V_{n}$. We have
$\mathfrak{hc}_{B}^{P_{k}}\iota_{P_{k}}({}_{k})\simeq\mathfrak{J}_{\boldsymbol{\omega}_{k}}=\mathfrak{J}_{1}\star\dots\star\mathfrak{J}_{k},$
by [BT22, Lemma 6.2.4]. We also have a natural unit morphism
[…]
@@
$\mathfrak{J}_{{\boldsymbol{\omega}}_{k}}$$\hat{\zeta}(\wedge^{k}\mathbf{V}_{n})$$\mathfrak{J}_{{\boldsymbol{\omega}}_{k}}$$\hat{\zeta}(\wedge^{k}\mathbf{V}_{n})$$\scriptstyle{w_{{\boldsymbol{\omega}}_{k}}}$$\scriptstyle{{\boldsymbol{\omega}}_{k}}$$\scriptstyle{E_{k}}$$\scriptstyle{w_{{\boldsymbol{\omega}}_{k}}}$
Here $w_{{\boldsymbol{\omega}}_{k}}=\mathfrak{hc}_{B}^{P_{k}}(r_{k}),E_{k}=\mathfrak{hc}(\varepsilon_{k})$, and ${\boldsymbol{\omega}}_{k}$ comes from the monodromy action on $\mathfrak{J}_{{\boldsymbol{\omega}}_{k}}$.
We define $w_{{\boldsymbol{\omega}}_{k}}$ to be the highest weight arrows for the images of fundamental representations $\wedge^{k}\mathbf{V}_{n}$. For a dominant weight $\boldsymbol{\lambda}=\sum a_{i}{\boldsymbol{\omega}}_{i},$ we must have, from the condition (d) of Proposition 4.1.2, $w_{\lambda}=\pi_{\boldsymbol{\lambda}}(w_{{\boldsymbol{\omega}}_{1}}^{\star a_{1}}\star\dots\star w_{\boldsymbol{\omega}_{n}}^{\star a_{n}})$, where $\pi_{\boldsymbol{\lambda}}$ stands for the projection $\pi_{\boldsymbol{\lambda}}:\hat{\zeta}(\otimes\mathbf{V}_{\boldsymbol{\omega}_{k}}^{\otimes a_{i}})\to\hat{\zeta}(\mathbf{V}_{\boldsymbol{\lambda}})$.
@@
The morphisms $E_{{\boldsymbol{\omega}}_{k}}$ obtained from $E_{1}$ by imposing condition (b) are equal to the morphisms $E_{k}$ defined above.
@@
$\operatorname{Ind}^{G}_{P_{1}}\operatorname{Res}^{G}_{P_{1}}(-)\to\mathcal{S}_{n}\star(-),$
@@
$\left(\operatorname{Ind}^{G}_{P_{1}}\operatorname{Res}^{G}_{P_{1}}\right)^{\circ k}(-)\to\mathcal{S}_{n}^{\star k}\star(-).$
By the proof of Proposition 3.4.2, it restricts to an isomorphism on the $S_{k}$-isotypic component corresponding to the sign representation. On the right-hand side this isotypic component is identified with the convolution $\hat{\zeta}(\wedge^{k}\mathbf{V}_{n})\star(-)$. On the left-hand side, the $S_{k}$-action can be described explicitly as follows. If a character sheaf $\mathcal{F}$ is given by the IC-extension of the local system corresponding to the $A_{n}$-module $M$, then $\operatorname{Ind}^{G}_{P}\operatorname{Res}^{G}_{P}\mathcal{F}$ is given by the IC-extension of the local system corresponding to a $A_{n}$-module $\mathbf{K}[S_{n}]\otimes_{\mathbf{K}[S_{n-1}]}M$. Choosing an isomorphism of $S_{n}$-modules
@@
and comparing the actions under this identification, we get that the action of $E_{{\boldsymbol{\omega}}_{k}}$ on the corresponding isotypic component is the same as given by $E_{k}$. ∎
4.3  Filtration by Jucys–Murphy sheaves.